YEAR 1 (R-Billion) YEAR 2 (R-Billion)

Investment 200 220



Saving 180 190

Export 100 110

Imports 120 140

Government Expenditure 150 160

Taxation 150 160

Equilibrium National Income 1 800 2 000



(a) Calculate the value of the multiplier for this economy.

(b) Should the full employment level of income be R-B2 255, by how much should government change its spending to reach this level of income during the next year given the multiplier has not changed.

(c) Evaluate whether or not this policy approach is effective in real life in achieving the desired level of GDP.
1
Expert's answer
2021-07-26T17:57:02-0400

(a)

The value of the multiplier is given by dividing the equilibrium national income by the government spending as follows:

"=\frac{1800}{150}"

"=12"

The value of the multiplier is "12\times".

(b)Given:

Level of income=2255

Value of multiplier =12

"Government Spending=\frac{2255}{12}"

"=187.9"

Government spending should be increased to 187.9 in order to reach the given level of income R-B 2255.

(c)

This policy is effective in achieving the desired level of GDP because it involves focusses mainly on the relationship between equilibrium national level of income and government expenditure.